Maggie has been collecting nickels and dimes for 2 weeks, and she has a total of 5.15. If she has 7 more nickles than she has dimes, how many of each coin does she have?
Answer by CubeyThePenguin(3113)

n = number of nickels
d = number of dimes

5n + 10d = 515
n = d + 7

Substitute n = d + 7.

5(d + 7) + 10d = 515
5d + 35 + 10d = 515
15d = 480
d = 32

There are 32 dimes and 39 nickels.